OpenStudy (anonymous):

Solve the following system of equations and show all work. y = −x2 + 4 y = 2x + 1

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I know the answers i just need help with how to get there

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Shalante any ideas??

OpenStudy (anonymous):

\[x^2\] is x^2 not x2

OpenStudy (anonymous):

yes, sorry it copy and pasted wrong

OpenStudy (anonymous):

if y=-x^2+4 and y=2x+1 then -x^2+4=2x+1 Know how to solve for x?

OpenStudy (anonymous):

not really lol

OpenStudy (anonymous):

if you do know how to solve for x, plug the x value to any of the equation to get y

OpenStudy (anonymous):

we don't know the x value

OpenStudy (anonymous):

\[-x^2+4=2x+1\]

OpenStudy (anonymous):

so solve and that is the solution?

OpenStudy (anonymous):

\[x^2+2x-3\] move it over

OpenStudy (anonymous):

yes you know how to solve that?

OpenStudy (anonymous):

there should be two solutions with both having (x,y) form

OpenStudy (anonymous):

what is the first step to solve that? lost again

OpenStudy (anonymous):

where do i go after the x^2+2x-3

OpenStudy (anonymous):

factor it ( )( )

OpenStudy (anonymous):

(x-1) (x+3) ? sorry I'm confused

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yea you are right.

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

dunno why it took me back to my old reply

OpenStudy (anonymous):

forgot to tell you its \[x^2+2x-3=0\]

OpenStudy (anonymous):

so quadratic ?

OpenStudy (anonymous):

that is what happens when you move all the numbers to the other side of an equation. After everything is gone, you a zero to it.

OpenStudy (anonymous):

yes i know

OpenStudy (anonymous):

so you get -3 after solving

OpenStudy (anonymous):

and 1

OpenStudy (anonymous):

so those are the x's what about the y's??

OpenStudy (anonymous):

Then plug both this x back into the first 2 equation using both x=-3 and x=1

OpenStudy (anonymous):

they should both equal right?

OpenStudy (anonymous):

yes they both get -5

OpenStudy (anonymous):

ANSWER #1 (x,y) plug in first x ANSWER #2 (x,y) plug in second x

OpenStudy (anonymous):

oops my orignial was wrong hold on

OpenStudy (anonymous):

so answer #1 is (-3,-5). you are correct on first one.

OpenStudy (anonymous):

I got everything under control! thanks so much!

OpenStudy (anonymous):

Yea, no problem.